A local travel agency donated the cruise, valued at 2000.00. The students sold 2852 raffle tickets at $5 per ticket. Kevin bought 1 ticket.
1.What is the probability that Kevin will win the spring break cruise?
2.What is the probability that Kevin will not win the cruise?
3. What is Kevin's expectation if he buys one ticket?
4. Interpret what the value calculated in no. 3 means about Kevin's expectation in participating in the raffle.
5. How does the value calculated in no. 3 compare to the amount Kevin paid for the ticket? Explain why the value in no. 3 is not exactly what Kevin paid for the ticket.

Answer by stanbon(75887)

A local travel agency donated the cruise, valued at 2000.00. The students sold 2852 raffle tickets at $5 per ticket. Kevin bought 1 ticket.
1.What is the probability that Kevin will win the spring break cruise?
P(win) = 1/2852
--------------------------------
2.What is the probability that Kevin will not win the cruise?
P(not win) = 2851/2852
--------------------------------
3. What is Kevin's expectation if he buys one ticket?
E(x) = (1/2852)(1995) + (2851/2852)(-5) = -$4.30
--------------------------------------
4. Interpret what the value calculated in no. 3 means about Kevin's expectation in participating in the raffle.
He can expect to lose $4.30 on each ticket he purcases.
---------------------------------------
5. How does the value calculated in no. 3 compare to the amount Kevin paid for the ticket?
Obvious
---------------------------------------
Explain why the value in no. 3 is not exactly what Kevin paid for the ticket.
Because he has a slight chance of winning and a large chance of losing.
----------
Cheers,
Stan